# Toward a Multi-Echelon Cyber Warfare Theory: A Meta-Game-Theoretic Paradigm for Defense and Dominance

### v1

@@ 4.3 The Meta-Game Warfare Equilibrium
In other words, a warfare equilibrium is stable under mutual recursion: no level can deviate without requiring re-adjustments in all others. Cyber conflict is therefore best understood not as a sequence of isolated games, but as an interdependent meta game, where equilibria at all levels must remain mutually consistent chen2019control ; zhu2015game ; huang2020dynamic . Any perturbation, whether a policy change at the top or a breakthrough in technology at the bottom, propagates through the hierarchy, transforming outcomes across every echelon.
@@
Figure 2: Recursive dependencies between equilibria at different levels of cyber warfare. Forward arrows (blue, solid) indicate constraints passed downward; backward arrows (red, dashed) capture feedback upward without overlapping. The technical equilibrium $E_{\text{technical}}=\theta^{*}$ influences all levels directly. A warfare equilibrium emerges only when forward and backward consistencies are both satisfied.
@@
Information structures and knowledge asymmetries define effective action sets and payoff functions li2022role ; yuksel2024stochastic ; bacsar2011prices . In some cases, more information improves strategic positioning, while in others it increases costs or generates uncertainty (e.g., information overload or deception). Thus, the equilibrium of warfare highlights that information can be both an advantage and a liability, depending on how it reshapes $E_{\text{operational}}$ or $E_{\text{strategic}}$.
@@
The recursive structure explains the nonlinear relationship between resources and advantage. An additional budget $B^{*}$ at the policy level does not automatically translate into dominance; rather, effectiveness depends on allocation strategies $r_{d}(\lambda)$ at the strategic level and adaptive tactics downstream. This captures the paradoxical logic of cyber conflict: more poorly deployed resources may yield little or no advantage.
@@
Finally, the concept of warfare equilibrium provides a foundation for designing resilience li2024symbiotic ; zhu2020crosslayer . Quick re-adjustment to perturbations, whether through tactical adaptation, operational redundancy, or strategic reserve allocation, ensures that the equilibrium can be restored without collapse. This recursive adaptability is what distinguishes fragile systems from resilient ones in cyber warfare.
@@ 4.4 Anamorphism and Catamorphism Across Echelons
To establish a rigorous framework, we formalize cross-echelon interactions through anamorphic and catamorphic mappings. These mappings capture the intuitive notions of zoom-in and zoom-out operations: anamorphisms expand coarse-grained decisions at a higher echelon into finer-grained actions at a lower echelon, while catamorphisms aggregate lower-level outcomes back into higher-level evaluations. Beyond their descriptive role, these mappings can be composed into hylomorphisms, which ensure consistency of the equilibrium across levels by requiring that the forward unfolding and the backward folding stabilize at a fixed point. This abstraction not only parallels constructions in category theory awodey2010category and program semantics meijer1991functional but also provides an analytical tool to reason about multi-level games and cyber-physical decision hierarchies ge2024mega ; zhu2015game ; chen2019control .
@@
$\mathsf{ana}_{O\to T}:\ \mathcal{A}_{d}\times\mathcal{A}_{a}\;\longrightarrow\;\Xi_{d}\times\Xi_{a},\qquad(a_{d}^{k},a_{a}^{k})\;\mapsto\;(\xi_{d},\xi_{a}),$
where $\mathcal{A}_{d}$ and $\mathcal{A}_{a}$ denote the defender’s and attacker’s operational action sets, while $\Xi_{d}$ and $\Xi_{a}$ denote their feasible sets of tactical sequences. This mapping represents the zoom-in process that refines a coarse operational choice into detailed tactical maneuvers.
@@
  • Strategy–Operation hylomorphism: $\mathsf{hylo}_{S\leftrightarrow O}\;\triangleq\;\mathsf{cata}_{O\to S}\circ\mathsf{ana}_{S\to O},$ which requires that strategic resource allocations, once expanded into operational policies and then aggregated back, remain aligned with the intended strategic evaluations.
  • Operation–Tactic hylomorphism: $\mathsf{hylo}_{O\leftrightarrow T}\;\triangleq\;\mathsf{cata}_{T\to O}\circ\mathsf{ana}_{O\to T},$ which guarantees that operational directives, when unfolded into tactical maneuvers and folded back, yield operational outcomes consistent with the original directives.
  • Policy–Strategy hylomorphism: $\mathsf{hylo}_{P\leftrightarrow S}\;\triangleq\;\mathsf{cata}_{S\to P}\circ\mathsf{ana}_{P\to S},$ which ensures that policy-level priorities (weights and budgets) are faithfully reflected in the equilibrium consequences observed at the strategic echelon.
Here, consistency means that a state is preserved under the round-trip mapping: if $x$ denotes a state at echelon $X$, then
$x=\mathsf{hylo}_{X\leftrightarrow Y}(x).$
In other words, unfolding the state into a lower echelon and folding it back again does not change the state. Such a fixed point indicates that decisions and outcomes are mutually coherent across the two levels, with no residual misalignment.
@@ 4.5 Warfare Equilibrium as a Lens on Stability, Dominance, and Winning
The equilibrium concept provides a unifying framework for analyzing warfare outcomes across multiple echelons. By definition, an equilibrium captures a state in which neither side can unilaterally improve its payoff given the actions of the other. Once embedded into our multi-echelon structure through anamorphisms, catamorphisms, and hylomorphisms, equilibria serve as a natural tool for addressing three fundamental notions in warfare: the stability of conflict dynamics, the dominance of one side over the other, and the winning of the war. Each builds on the other, forming a progression from robustness, to superiority, to decisive success.
@@
This happens for several reasons. First, strategy often involves trade-offs: resources may be conserved in one theater to achieve a decisive advantage in another. Second, local setbacks can be part of deception, drawing the adversary into costly commitments that ultimately reduce their overall payoff. Finally, what defines victory is not the outcome of isolated engagements, but whether policy objectives are achieved, such as deterrence, survival, or coalition cohesion. In equilibrium terms, the defender still “wins the war” if the global hylomorphic mapping preserves consistency and produces a policy-level state in which the payoff of the defender meets or exceeds its required threshold, while the payoff of the attacker remains at or below its corresponding threshold.
Thus, losing a battle at the tactical or operational echelon does not equate to losing the war. The essential point is that war is won at the level of the global equilibrium, where tactical and operational outcomes are subordinated to strategic coherence and policy success.
@@ 5 Cyber Warfare Taxonomy
From a perspective of capabilities, asymmetric conflicts, where weaker actors exploit low-cost attacks against stronger adversaries, are well captured by attacker–defender security games, with unbalanced action sets or unequal costs. Symmetric conflicts, where both sides have comparable capabilities, lend themselves to zero-sum matrix games, where $u_{d}=-u_{a}$. Escalatory campaigns are naturally modeled as repeated or stochastic games, in which state transitions $s^{t+1}$ reflect cycles of action and counteraction.
A taxonomy by actors determines the set of players $N$. Nation-states, proxies, non-state groups, and insiders all enter with different information, resources, and objectives. State–state interactions are often modeled as strategic-form or signaling games, while insider threats require principal–agent models to capture hidden actions and incentive misalignments within organizations.
@@
Finally, classification by methods aligns with the appropriate game structures under uncertainty. Sabotage and disruption are modeled with stochastic or differential games to capture cascading effects. Espionage fits naturally into a Bayesian game, with hidden types and incomplete information. Disinformation and propaganda are framed as signaling or information design games, where one player manipulates beliefs to influence the other’s response.
@@
Actors States, Proxies, Non-State Groups, Insiders Strategic-form or signaling games; Principal–agent models (hidden actions); Bayesian games (incomplete information).
@@ 6.1 Current State
The cyber conflict between China and Taiwan has escalated alongside broader geopolitical tensions. Episodes around politically sensitive events, such as high-level visits or elections, often coincide with surges in cyber operations. These include distributed denial-of-service (DDoS) attacks, misinformation campaigns, and targeted infrastructure probing. For example, during Nancy Pelosi’s visit to Taiwan in August 2022, cyberattacks hijacked digital billboards to display anti-Pelosi messages, revealing vulnerabilities in Taiwan’s information infrastructure. Similarly, in April 2023, when Taiwanese President Tsai Ing-wen met U.S. House Speaker Kevin McCarthy, China coupled military maneuvers with coordinated cyberattacks, signaling a synchronized dual-pronged strategy. Other incidents, such as the severing of undersea cables to Taiwan’s Matsu Islands in February 2023, underscored the fragility of subsea infrastructure and the need for redundant communication paths. Taiwan’s response, ranging from emergency microwave relays to LEO satellite deployment, demonstrates a growing emphasis on institutional resilience. The creation of the Ministry of Digital Affairs (MODA) and agencies such as the National Institute of Cyber Security highlights the island’s long-term strategy of integrating cybersecurity with governance and innovation.
@@ 6.2 RedCyber Warfare: Coordinated Cyber Campaign on Taiwan
As visualized in Fig. 3, RedCyber unfolds as a stage-structured game that pushes from policy to reconnaissance, disruption, escalation, sustained assault, and strategic messaging, with solid arrows showing top–down propagation of objectives/budgets and dashed arrows showing bottom–up feedback from tactical/technical outcomes to higher echelons. At each stage $k$, China’s actions $a_{a}^{k}$ and Taiwan’s defenses $a_{d}^{k}$ co-evolve via
@@
Months before a visible crisis, Beijing’s doctrine blends intrusion, economic pressure, and information ops; Taipei counters by setting weights and budgets $(w,B)$, committing to deception portfolios $\lambda$, and activating alliances that provide surge capacity. In Fig. 3, this is the Policy node (k=0k{=}0), which pushes a feasible game downstream by constraining the resources and timing available to the Strategic and Operational nodes. Formally, the policy equilibrium $E_{\mathrm{policy}}$ parameterizes $\mathcal{G}_{\mathrm{strategic}}$ and the budget sets therein, consistent with cross-echelon meta-game design yang2025multires ; huang2020dynamic ; chen2019dynamic . Empirically, Taiwan’s establishment of the Ministry of Digital Affairs (MODA) and associated resilience programs provides an institutional anchor for $(w,B,\lambda)$ and $\theta$ updates moda-site .
@@
APT teams exploit zero-days in edge devices and conduct tailored phishing of municipal IT staff; sockpuppets seed test narratives on social platforms to probe which themes resonate. Taiwan plants decoy credentials and canary tokens to induce observable errors and generate micro-evidence for attribution without revealing coverage. This corresponds to asymmetric-information/signaling-with-evidence games (the Recon node in Fig. 3), where the attacker learns hidden types and the defender manipulates observability pawlick2018leaky ; zhang2018hypothesis ; huang2019adaptive ; huang2020farsighted ; huang2021duplicity . Episodes around sensitive politics illustrate the timing: during Speaker Pelosi’s August 2022 visit, Taiwan experienced surges in DDoS and screen hijacks of public displays reuters-pelosi-2022 ; abc-pelosi-2022 ; cyberscoop-2022 .
@@
With footholds established, malware on industrial control systems (ICS) triggers localized outages while DDoS floods emergency portals; fabricated outage maps amplify panic. Taiwan’s counter is to island microgrids, shift traffic to prepositioned static mirrors and CDN, and create emergency communications (ERCV / LEO terminals) to maintain command and control. These interactions match sequential attack–defense games on the CPS (the Disruption node), with hybrid/stochastic models that capture cascades and rapid response miao2018hybrid ; zhao2020semiNP ; chen2019dynamic . The link to practice is visible in the response to the Hualien earthquake of Taiwan on April 3, 2024, where MODA deployed the first Mobile Emergency Network Vehicle and later airlifted LEO satellite terminals to restore connectivity, a concrete increase in $\theta$ that steepens attacker cost curves in $T(\cdot)$ moda-ercv-2024a ; moda-ercv-2024b .
@@
As elections approach, credential-stuffing and spoofed “mirror” results appear; deepfakes and forged documents attempt to delegitimize candidates; semiconductor and transport disruptions compound shocks. Taiwan activates real-time audits, paper backups, golden-image rollbacks, and transparent livestreams. The Escalation node is well captured as a Stackelberg game: the attacker leads with politically/economically coupled moves; the defender responds under resource constraints liu2025stackrisk . Information-design tools formalize the overt/covert counter-messaging trade-offs li2023price . In practice, cyber and kinetic signaling has been tightly coupled around high-visibility diplomacy and exercises icg-2023 . Cloud measurement shows Taiwan-bound DDoS spiking around election periods, aligning with timing games that exploit attention cycles cloudflare-q4-2023 ; cloudflare-q1-2023 .
@@
The tempo settles into weekly breach dumps, payment hiccups, and refresh spearphishing to re-seed access. Taiwan rotates keys (including PQC pilots), segments high-value data, automates fraud detection, and uses insurance/contracting to share risk. This is the Sustained Assault node, modeled as a repeated game in which resilience investments and incentive instruments depress the attacker’s marginal returns zhu2020crosslayer ; zhang2019insurance ; liu2025stackrisk . Physical infrastructure remains a live risk channel; repeated damage to undersea cables near the Matsu islands has forced backup communication and hardened procedures, illustrating how exogenous shocks perturb $T(\cdot)$ and the feasible action sets reuters-matsu-2025 ; ap-matsu-2023 .
@@
Finally, partial restoration of services is overlaid with inevitability narratives; embassy spoofs and “relief phishing” siphon credentials. Taiwan responds with cryptographically signed fact-checks, international amplification of verified messages, and rapid takedown of coordinated inauthentic behavior. The Strategic Messaging node is a signaling (information-design) game in which credibility, transparency, and commitment shape equilibrium beliefs and blunt coercion li2023price ; liu2023dim .
@@
Throughout, the deception index $\lambda$ and technology lever $\theta$ (annotated above the timeline in Fig. 3) move the frontier of feasible actions and detection latencies. In equilibrium terms, $E_{\text{policy}}$ sets budgets, alliances, and doctrine that parameterize $\mathcal{G}_{\text{strategic}}$; $E_{\text{strategic}}$ allocates resources/timing that constrain $\mathcal{G}_{\text{operational}}$; operational outcomes filter feasible tactics and belief states; and tactical/technical outcomes feed back (dashed arrows) to shift higher-level equilibria by changing costs and beliefs. Advances in $\theta$, notably LLM/agentic-AI defenses for detection, reasoning, and counter-deception, have been modeled directly as LLM-Nash/Stackelberg games and deception timing games; these mechanisms compress attacker windows and reshape upstream plans zhu2025llmnash ; zhu2025llmstack ; li2024symbiotic ; yang2025timing . In short, RedCyber behaves as meta-games: The equilibrium of each node depends on and perturbs the rest, with real-world incidents providing clear calibration points for $T(\cdot)$, costs and beliefs.
Jan–Mar 2024Policy & Preparation (k=0k{=}0)Goals, budgets, alliancesApr–Sep 2024Reconnaissance ($k_{1}$)Asymmetric info; $\lambda$ decoys/canariesOct–Dec 2024Initial Disruption ($k_{2}$)CPS/DDoS; $\theta$ shifts $T(\cdot)$Jan–Mar 2025Escalation ($k_{3}$)Leader–follower responsesApr 2025–ongoingSustained OperationsRepeated play; incentives/insurance
Figure 3: Stage-structured timeline of the RedCyber campaign against Taiwan. Solid arrows represent top–down propagation of goals, budgets, and doctrine from the policy echelon to downstream games, while dashed arrows capture bottom–up feedback of tactical and technical outcomes to higher levels. Each node illustrates a phase of the campaign, policy preparation, reconnaissance, disruption, escalation, and sustained operations, with the deception index $\lambda$ and technology lever $\theta$ annotated as cross-cutting parameters that shift detection latencies and feasible action sets in $T(\cdot)$.

### v2

@@ 4.3 The Meta-Game Warfare Equilibrium
In other words, a warfare equilibrium is stable under mutual recursion: no level can deviate without requiring re-adjustments in all others. Cyber conflict is therefore best understood not as a sequence of isolated games, but as an interdependent meta game, where equilibria at all levels must remain mutually consistent Chen and Zhu (2019b); Zhu and Başar (2015); Huang et al. (2020). Any perturbation, whether a policy change at the top or a breakthrough in technology at the bottom, propagates through the hierarchy, transforming outcomes across every echelon.
@@
Figure 2: Recursive interdependencies among equilibria across policy, strategic, operational, and tactical levels of cyber warfare. Forward (blue, solid) arrows represent constraints and decisions propagated downward across echelons, while backward (red, dashed) arrows capture upward feedback effects. The technical equilibrium $E_{\text{technical}}=\theta^{*}$ influences all levels directly. A coherent warfare equilibrium emerges only when both forward constraints and backward feedback are jointly satisfied, ensuring cross-echelon consistency.
@@
Information structures and knowledge asymmetries define effective action sets and payoff functions Li et al. (2022); Yüksel and Başar (2024); Başar and Zhu (2011). In some cases, more information improves strategic positioning, while in others it increases costs or generates uncertainty (e.g., information overload or deception). Thus, the equilibrium of warfare highlights that information can be both an advantage and a liability, depending on how it reshapes $E_{\text{operational}}$ or $E_{\text{strategic}}$.
@@
The recursive structure explains the nonlinear relationship between resources and advantage. An additional budget $B^{*}$ at the policy level does not automatically translate into dominance; rather, effectiveness depends on allocation strategies $r_{d}(\lambda)$ at the strategic level and adaptive operations downstream. This captures the paradoxical logic of cyber conflict: more poorly deployed resources may yield little or no advantage.
@@
Finally, the concept of warfare equilibrium provides a foundation for designing resilience Li and Zhu (2025); Zhu and Xu (2020). Quick re-adjustment to perturbations, whether through tactical adaptation, operational redundancy, or strategic reserve allocation, ensures that the equilibrium can be restored without collapse. This recursive adaptability is what distinguishes fragile systems from resilient ones in cyber warfare.
@@ 4.4 Anamorphism and Catamorphism Across Echelons
To establish a rigorous framework, we formalize cross-echelon interactions through anamorphic and catamorphic mappings. These mappings capture the intuitive notions of zoom-in and zoom-out operations: anamorphisms expand coarse-grained decisions at a higher echelon into finer-grained actions at a lower echelon, while catamorphisms aggregate lower-level outcomes back into higher-level evaluations. Beyond their descriptive role, these mappings can be composed into hylomorphisms, which ensure consistency of the equilibrium across levels by requiring that the forward unfolding and the backward folding stabilize at a fixed point. This abstraction not only parallels constructions in category theory Awodey (2010) and program semantics Meijer et al. (1991) but also provides an analytical tool to reason about multi-level games and cyber-physical decision hierarchies Ge and Zhu (2024); Zhu and Başar (2015); Chen and Zhu (2019b).
@@
$\mathsf{ana}_{O\to T}:\ A_{d}\times A_{a}\;\longrightarrow\;\Xi_{d}\times\Xi_{a},\qquad(a_{d}^{k},a_{a}^{k})\;\mapsto\;(\xi_{d},\xi_{a}),$
where $A_{d}$ and $A_{a}$ denote the defender’s and attacker’s operational action sets, while $\Xi_{d}$ and $\Xi_{a}$ denote their feasible sets of tactical sequences. This mapping represents the zoom-in process that refines a coarse operational choice into detailed tactical maneuvers.
@@
  • Strategy-Operation hylomorphism: $\mathsf{hylo}_{S\leftrightarrow O}\;\triangleq\;\mathsf{cata}_{O\to S}\circ\mathsf{ana}_{S\to O},$ which requires that strategic resource allocations, once expanded into operational policies and then aggregated back, remain aligned with the intended strategic evaluations.
  • Operation-Tactic hylomorphism: $\mathsf{hylo}_{O\leftrightarrow T}\;\triangleq\;\mathsf{cata}_{T\to O}\circ\mathsf{ana}_{O\to T},$ which guarantees that operational directives, when unfolded into tactical maneuvers and folded back, yield operational outcomes consistent with the original directives.
  • Policy-Strategy hylomorphism: $\mathsf{hylo}_{P\leftrightarrow S}\;\triangleq\;\mathsf{cata}_{S\to P}\circ\mathsf{ana}_{P\to S},$ which ensures that policy-level priorities (weights and budgets) are faithfully reflected in the equilibrium consequences observed at the strategic echelon.
Here, consistency means that a state is preserved under the round-trip mapping: if $x$ denotes a state at echelon $X$, then $x=\mathsf{hylo}_{X\leftrightarrow Y}(x).$ In other words, unfolding the state into a lower echelon and folding it back again does not change the state. Such a fixed point indicates that decisions and outcomes are mutually coherent across the two levels, with no residual misalignment.
@@ 4.5 Warfare Equilibrium as a Lens on Stability, Dominance, and Winning
The equilibrium concept provides a unifying framework for analyzing warfare outcomes across multiple echelons. By definition, an equilibrium captures a state in which neither side can unilaterally improve its payoff given the actions of the other. Once embedded into our multi-echelon structure through anamorphisms, catamorphisms, and hylomorphisms, equilibria serve as a natural tool for addressing three fundamental notions in warfare: the stability of conflict dynamics, the dominance of one side over the other, and the winning of the war. Each builds on the other, forming a progression from robustness to superiority to decisive success.
@@
This happens for several reasons. First, strategy often involves trade-offs: resources may be conserved in one theater to achieve a decisive advantage in another. Second, local setbacks can be part of deception, drawing the adversary into costly commitments that ultimately reduce their overall payoff. Finally, what defines victory is not the outcome of isolated engagements, but whether policy objectives are achieved, such as deterrence, survival, or coalition cohesion. In equilibrium terms, the defender still “wins the war” if the global hylomorphic mapping preserves consistency and produces a policy-level state in which the defender’s payoff meets or exceeds its required threshold, while the attacker’s payoff remains at or below its corresponding threshold.
Thus, losing a battle at the tactical or operational echelon does not equate to losing the war. The essential point is that war is won at the global equilibrium level, where tactical and operational outcomes are subordinated to strategic coherence and policy success.
@@ 5 Cyber Warfare Taxonomy
From a perspective of capabilities, asymmetric conflicts, where weaker actors exploit low-cost attacks against stronger adversaries, are well captured by attacker-defender security games, with unbalanced action sets or unequal costs. Symmetric conflicts, where both sides have comparable capabilities, lend themselves to zero-sum matrix games, where $u_{d}=-u_{a}$. Escalatory campaigns are naturally modeled as repeated or stochastic games, in which state transitions $s^{t+1}$ reflect cycles of action and counteraction.
A taxonomy by actors determines the set of players $N$. Nation-states, proxies, non-state groups, and insiders all enter with different information, resources, and objectives. State-state interactions are often modeled as strategic-form or signaling games, while insider threats require principal-agent models to capture hidden actions and incentive misalignments within organizations Zhang and Zhu (2019); Zhu (2025b); Zhu and Başar (2025).
@@
Finally, classification by methods aligns with the appropriate game structures under uncertainty. Sabotage and disruption are modeled with stochastic or differential games to capture cascading effects. Espionage fits naturally into a Bayesian game, with hidden types and incomplete information. Disinformation and propaganda are framed as signaling or information design games Yang et al. (2023); Zhang and Zhu (2021b, a), where one player manipulates beliefs to influence the other’s response.
@@
Actors States, Proxies, Non-State Groups, Insiders Strategic-form or signaling games; Bayesian games (incomplete information); Principal-agent models (hidden actions).
@@ 6.1 Current State
The cyber conflict between China and Taiwan has escalated alongside broader geopolitical tensions and increasingly reflects characteristics of hybrid warfare Solmaz (2024). Episodes around politically sensitive events, such as high-level visits or elections, often coincide with surges in cyber operations. These include DDoS attacks, misinformation campaigns, cyber-espionage, and targeted infrastructure probing, consistent with broader global threat trends documented in recent cybersecurity assessments European Union Agency for Cybersecurity (ENISA) (2025); Microsoft Threat Intelligence (2023).
For example, during Nancy Pelosi’s visit to Taiwan in August 2022, cyberattacks hijacked digital billboards to display anti-Pelosi messages, revealing vulnerabilities in Taiwan’s public information infrastructure. Similarly, in April 2023, when Taiwanese President Tsai Ing-wen met U.S. House Speaker Kevin McCarthy, China coupled military maneuvers with coordinated cyber activities, signaling an integrated cross-domain pressure strategy Solmaz (2024).
Other incidents, such as the severing of undersea cables to Taiwan’s Matsu Islands in February 2023, underscored the fragility of communication infrastructure and the strategic importance of redundancy. In response, Taiwan has accelerated resilience initiatives, including emergency microwave relays and the development of a low-Earth-orbit (LEO) satellite communications ecosystem Wang (2026); SpaceX (2023). These efforts reflect a broader shift toward diversified and space-enabled connectivity to mitigate geopolitical risk.
Institutionally, Taiwan has strengthened its governance framework through the establishment of the Ministry of Digital Affairs (MODA) and the expansion of agencies such as the National Institute of Cyber Security Government of Taiwan (2022); National Institute of Cyber Security, Taiwan (2023). These reforms align with the objectives outlined in Taiwan’s National Cyber Security Program (2021–2024), which emphasizes infrastructure protection, public–private coordination, and strategic resilience planning National Centre for Cyber Security Technology (2021). Collectively, these developments signal a transition from reactive cybersecurity measures toward a comprehensive resilience-oriented national strategy.
@@ 6.2 RedCyber Warfare: Coordinated Cyber Campaign on Taiwan
As visualized in Fig. 3, RedCyber unfolds as a stage-structured game that pushes from policy to reconnaissance, disruption, escalation, sustained assault, and strategic messaging, with solid arrows showing top-down propagation of objectives/budgets and dashed arrows showing bottom-up feedback from tactical/technical outcomes to higher echelons. At each stage $k$, China’s actions $a_{a}^{k}$ and Taiwan’s defenses $a_{d}^{k}$ co-evolve via
@@
Months before a visible crisis, Beijing’s doctrine blends intrusion, economic pressure, and information ops; Taipei counters by setting weights and budgets $(w,B)$, committing to deception portfolios $\lambda$, and activating alliances that provide surge capacity. In Fig. 3, this is the Policy node (k=0k{=}0), which pushes a feasible game downstream by constraining the resources and timing available to the Strategic and Operational nodes. Formally, the policy equilibrium $E_{\mathrm{policy}}$ parameterizes $\mathcal{G}_{\mathrm{strategic}}$ and the budget sets therein, consistent with cross-echelon meta-game design Yang and Zhu (2026); Huang et al. (2020); Chen and Zhu (2019a). Empirically, Taiwan’s establishment of MODA and associated resilience programs provides an institutional anchor for $(w,B,\lambda)$ and $\theta$ updates Ministry of Digital Affairs, Taiwan (2022).
@@
APT teams exploit zero-days in edge devices and conduct tailored phishing of municipal IT staff; sock puppets seed test narratives on social platforms to probe which themes resonate. Defenders plant decoy credentials and canary tokens to induce observable errors and generate micro-evidence for attribution without revealing coverage. This corresponds to asymmetric-information/signaling-with-evidence games (the Recon node in Fig. 3), where the attacker learns hidden types and the defender manipulates observability Pawlick et al. (2018a); Zhang and Zhu (2018); Huang and Zhu (2019a, 2020, 2021b). Episodes around sensitive politics illustrate the timing: during Speaker Pelosi’s August 2022 visit, Taiwan experienced surges in DDoS and screen hijacks of public displays Reuters (2022); Australian Broadcasting Corporation (2022); Vicens (2022).
@@
With footholds established, malware on industrial control systems (ICS) triggers localized outages while DDoS floods emergency portals; fabricated outage maps amplify panic. Taiwan’s counter is to island microgrids, shift traffic to prepositioned static mirrors and CDN, and create emergency communications (ERCV / LEO terminals) to maintain command and control. These interactions match sequential attack-defense games on the CPS (the Disruption node), with hybrid/stochastic models that capture cascades and rapid response Miao et al. (2018); Zhao et al. (2020); Chen and Zhu (2019a). The link to practice is visible in the response to the Hualien earthquake of Taiwan on April 3, 2026, where MODA deployed the first Mobile Emergency Network Vehicle and later airlifted LEO satellite terminals to restore connectivity, a concrete increase in $\theta$ that steepens attacker cost curves in $T(\cdot)$ Ministry of Digital Affairs, Taiwan (2024b, a).
@@
As elections approach, credential-stuffing and spoofed “mirror” results appear; deepfakes and forged documents attempt to delegitimize candidates; semiconductor and transport disruptions compound shocks. Taiwan activates real-time audits, paper backups, golden-image rollbacks, and transparent livestreams. The Escalation node is well captured as a Stackelberg game: the attacker leads with politically/economically coupled moves; the defender responds under resource constraints Liu and Zhu (2025). Information-design tools formalize the overt/covert counter-messaging trade-offs Li and Zhu (2023). In practice, cyber and kinetic signaling have been tightly coupled around high-visibility diplomacy and exercises International Crisis Group (2023). Cloud measurement shows Taiwan-bound DDoS spiking around election periods, aligning with timing games that exploit attention cycles Cloudflare (2024, 2023).
@@
The tempo settles into weekly breach dumps, payment hiccups, and refresh spearphishing to re-seed access. Taiwan rotates keys (including PQC pilots), segments high-value data, automates fraud detection, and uses insurance/contracting to share risk. This is the Sustained Assault node, can be modeled as a repeated game in which resilience investments and incentive instruments depress the attacker’s marginal returns Zhu and Xu (2020); Zhang and Zhu (2019); Liu and Zhu (2025). Physical infrastructure remains a live risk channel; repeated damage to undersea cables near the Matsu islands has forced backup communication and hardened procedures, illustrating how exogenous shocks perturb $T(\cdot)$ and the feasible action sets Reuters (2025); Associated Press (2023).
@@
Finally, partial restoration of services is overlaid with inevitability narratives; embassy spoofs and “relief phishing” siphon credentials. Taiwan responds with cryptographically signed fact-checks, international amplification of verified messages, and rapid take-down of coordinated inauthentic behavior. The Strategic Messaging node is a signaling (information-design) game in which credibility, transparency, and commitment shape equilibrium beliefs and blunt coercion Li and Zhu (2023); Liu and Zhu (2023).
@@
Throughout, the deception index $\lambda$ and technology lever $\theta$ (annotated above the timeline in Fig. 3) move the frontier of feasible actions and detection latencies. In equilibrium terms, $E_{\text{policy}}$ sets budgets, alliances, and doctrine that parameterize $\mathcal{G}_{\text{strategic}}$; $E_{\text{strategic}}$ allocates resources/timing that constrain $\mathcal{G}_{\text{operational}}$; operational outcomes filter feasible tactics and belief states; and tactical/technical outcomes feed back (dashed arrows) to shift higher-level equilibria by changing costs and beliefs. Advances in $\theta$, notably LLM/agentic-AI defenses for detection, reasoning, and counter-deception, have been modeled directly as LLM-Nash/Stackelberg games and deception timing games; these mechanisms compress attacker windows and reshape upstream plans Zhu (2025d, c); Li and Zhu (2025); Yang and Zhu (2025b). In short, RedCyber behaves as meta-games: The equilibrium of each node depends on and perturbs the rest, with real-world incidents providing clear calibration points for $T(\cdot)$, costs and beliefs.
Jan–Mar 2026Policy & Preparation (k=0k{=}0)Goals, budgets, alliancesApr–Sep 2026Reconnaissance ($k_{1}$)Asymmetric info; $\lambda$ decoys/canariesOct–Dec 2026Initial Disruption ($k_{2}$)CPS/DDoS; $\theta$ shifts $T(\cdot)$Jan–Mar 2027Escalation ($k_{3}$)Leader–follower responsesApr 2027–ongoingSustained OperationsRepeated play; incentives/insurance
Figure 3: Stage-structured timeline of the RedCyber campaign against Taiwan. Each node illustrates a phase of the campaign, policy preparation, reconnaissance, disruption, escalation, and sustained operations, with the deception index $\lambda$ and technology lever $\theta$ annotated as cross-cutting parameters that shift detection latencies and feasible action sets in $T(\cdot)$.